OpenStudy (anonymous):

rimshot has two tv sets. the area of the smaller screen is 40% of the area of the larger screen. if the larger screen measures 24 in. by 32 in, what is the area of the smaller screen?

OpenStudy (anonymous):

x/24=y/32 and xy=307. Solve the system.

OpenStudy (anonymous):

307.2 in.^2

OpenStudy (anonymous):

I'm of the sliderule generation; I rounded it off.

OpenStudy (anonymous):

Darn, misread problem. Sorry, made it a lot harder than it needed to be....

OpenStudy (anonymous):

Getting late, I guess....LOL